# Out-of-bag prediction balls for random forests in metric spaces

### v1

@@ Appendix D Connection of the HvMF and vMF distributions and normalizing constants
The Hyperboloid von–Mises Fisher (HvMF) distribution on $\mathbb{H}^{d}$, $d\geq 1$, with location $\boldsymbol{\mu}\in\mathbb{H}^{d}$ and concentration $\kappa>0$ was introduced by Jensen, (1981) as that induced by the density
@@
In the following proposition, we reapproach the derivations by (Jensen,, 1981, pp. 196–197) to clarify the connection of the HvMF and vMF distributions, and to detail the derivation and computation of the normalizing constant. The complexity of computing the normalizing constant for the HvMF is comparable to that of the vMF for even dimension $d$ and is slightly more direct for odd $d$, in the sense of simplifying the finite series
@@
for integers $m\geq 0$ (see Equation 8.467 in Zwillinger et al., (2014)) to the series
@@
Following Jensen, (1981), for $\boldsymbol{x}\in\mathbb{H}^{d}$ and $u\in\mathbb{R}$, define the hyperbolic-spherical coordinates:
@@
The group of hyperbolic transformations acts transitively on $\mathbb{H}^{d}$ (see Lemma 2(ii) of Jensen,, 1981), i.e., for any pair of mean vectors $\boldsymbol{\mu},\tilde{\boldsymbol{\mu}}\in\mathbb{H}^{d}$, there exists an action $\boldsymbol{A}\in\mathrm{SH}(d)$ such that $\tilde{\boldsymbol{\mu}}=\boldsymbol{A}\boldsymbol{\mu}$. The equality $c_{d}^{\mathrm{HvMF}}(\boldsymbol{\mu},\kappa)=c_{d}^{\mathrm{HvMF}}(\tilde{\boldsymbol{\mu}},\kappa)$ follows from
@@
The second equality in (37) holds because the Minkowski pseudo-inner product is invariant under hyperbolic transformations, i.e., $(\boldsymbol{x},\boldsymbol{y})=(\boldsymbol{A}\boldsymbol{x},\boldsymbol{A}\boldsymbol{y})$ for every $\boldsymbol{x},\boldsymbol{y}\in\mathbb{H}^{d}$ and every $\boldsymbol{A}\in\mathrm{SH}(d)$. The invariance follows from the identity $(\boldsymbol{x},\boldsymbol{y})=\boldsymbol{x}^{\top}\tilde{\boldsymbol{I}}_{d+1}\boldsymbol{y}$ for any $\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{d+1}$. For the first equality in (38), first observe that for every $\boldsymbol{A}\in\mathrm{SH}(d)$, its inverse $\boldsymbol{A}^{-1}\in\mathrm{SH}(d)$ is given by $\boldsymbol{A}^{-1}=\tilde{\boldsymbol{I}}_{d+1}\boldsymbol{A}^{\top}\tilde{\boldsymbol{I}}_{d+1}$. Now, we apply the substitution $\boldsymbol{x}=\boldsymbol{A}^{-1}\boldsymbol{z}$ (note that $\det(\boldsymbol{A}^{-1})=1$). Observe also that the domain of integration remains invariant since the mapping $\boldsymbol{x}\mapsto\boldsymbol{A}^{-1}\boldsymbol{x}$ is a bijection from $\mathbb{H}^{d}$ to itself. Surjectivity holds because for any $\boldsymbol{z}\in\mathbb{H}^{d}$, $\boldsymbol{z}=\boldsymbol{A}^{-1}(\boldsymbol{A}\boldsymbol{x})$, and $\boldsymbol{A}\boldsymbol{x}\in\mathbb{H}^{d}$ because the unit hyperboloid is invariant under hyperbolic transformations (Lemma 1(ii) of Jensen,, 1981). ∎

### v2

@@ Appendix D Connection of the HvMF and vMF distributions and normalizing constants
The Hyperboloid von–Mises Fisher (HvMF) distribution on $\mathbb{H}^{d}$, $d\geq 1$, with location $\boldsymbol{\mu}\in\mathbb{H}^{d}$ and concentration $\kappa>0$ was introduced by Jensen (1981) as that induced by the density
@@
In the following proposition, we reapproach the derivations by (Jensen, 1981, pp. 196–197) to clarify the connection of the HvMF and vMF distributions, and to detail the derivation and computation of the normalizing constant. The complexity of computing the normalizing constant for the HvMF is comparable to that of the vMF for odd dimension $d$ and is slightly more direct for even $d$, in the sense of simplifying the finite series
@@
for integers $m\geq 0$ (see Equation 8.467 in Zwillinger et al. (2014)) to the series
@@
Following Jensen (1981), for $\boldsymbol{x}\in\mathbb{H}^{d}$ and $u\in\mathbb{R}$, define the hyperbolic-spherical coordinates:
@@
The group of hyperbolic transformations acts transitively on $\mathbb{H}^{d}$ (see Lemma 2(ii) of Jensen, 1981), i.e., for any pair of mean vectors $\boldsymbol{\mu},\tilde{\boldsymbol{\mu}}\in\mathbb{H}^{d}$, there exists an action $\boldsymbol{A}\in\mathrm{SH}(d)$ such that $\tilde{\boldsymbol{\mu}}=\boldsymbol{A}\boldsymbol{\mu}$. The equality $c_{d}^{\mathrm{HvMF}}(\boldsymbol{\mu},\kappa)=c_{d}^{\mathrm{HvMF}}(\tilde{\boldsymbol{\mu}},\kappa)$ follows from
@@
The second equality in (37) holds because the Minkowski pseudo-inner product is invariant under hyperbolic transformations, i.e., $(\boldsymbol{x},\boldsymbol{y})=(\boldsymbol{A}\boldsymbol{x},\boldsymbol{A}\boldsymbol{y})$ for every $\boldsymbol{x},\boldsymbol{y}\in\mathbb{H}^{d}$ and every $\boldsymbol{A}\in\mathrm{SH}(d)$. The invariance follows from the identity $(\boldsymbol{x},\boldsymbol{y})=\boldsymbol{x}^{\top}\tilde{\boldsymbol{I}}_{d+1}\boldsymbol{y}$ for any $\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{d+1}$. For the first equality in (38), first observe that for every $\boldsymbol{A}\in\mathrm{SH}(d)$, its inverse $\boldsymbol{A}^{-1}\in\mathrm{SH}(d)$ is given by $\boldsymbol{A}^{-1}=\tilde{\boldsymbol{I}}_{d+1}\boldsymbol{A}^{\top}\tilde{\boldsymbol{I}}_{d+1}$. Now, we apply the substitution $\boldsymbol{x}=\boldsymbol{A}^{-1}\boldsymbol{z}$ (note that $\det(\boldsymbol{A}^{-1})=1$). Observe also that the domain of integration remains invariant since the mapping $\boldsymbol{x}\mapsto\boldsymbol{A}^{-1}\boldsymbol{x}$ is a bijection from $\mathbb{H}^{d}$ to itself. Surjectivity holds because for any $\boldsymbol{z}\in\mathbb{H}^{d}$, $\boldsymbol{z}=\boldsymbol{A}^{-1}(\boldsymbol{A}\boldsymbol{x})$, and $\boldsymbol{A}\boldsymbol{x}\in\mathbb{H}^{d}$ because the unit hyperboloid is invariant under hyperbolic transformations (Lemma 1(ii) of Jensen, 1981). ∎